Sales Mix and Break-Even Analysis Heyden Company has fixed costs of $773,480. The unit selling price, variable cost per unit, and contribution margin per unit for the company's two products follow: Product Selling Price Variable Cost per Unit Contribution Margin per Unit QQ $240 $160 $80 ZZ 300 160 140 The sales mix for Products QQ and ZZ is 30% and 70%, respectively. Determine the break-even point in units of QQ and ZZ. If required, round your answers to the nearest whole number. a. Product QQ units b. Product ZZ units
Answer #1

ANSWER:

Calculation of break even point in units:

Break even point in units= Fixed cost/ contribution per unit

Contribution per unit=80*0.30+140*0.70= $122

Break even point= 773480/122= 6340 units

a) Product QQ= 6340*0.30= 1902 units

b) Product ZZ= 6340*0.70=4438 units

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************